OpenStudy (anonymous):

Claude and Rosalinda are debating whether an equation from their homework assignment is an indentily .Claude says that some he has tried ten specific value for the variable and all of them worked ,it must be used as counterexamples Roselinda explain that specific values could only be use as counterexamples to prove that an equation is not an identity.Who is correct? .Explain your answer.

OpenStudy (anonymous):

I swear you just asked this, and even gave me a medal for it

OpenStudy (anonymous):

I did give you medal, are you delete answer? I can't find

OpenStudy (anonymous):

It seems you deleted the previous question, so I will answer it again. Roselinda is correct. Consider the following conjecture: 3q + 1 always results in an odd number. This is not the case. We can show some cases where this is true, however: 3x2 + 1 = 7, which is odd 3x4 + 1 = 13, which is odd, and we can continue in this way. However, this does NOT mean it is true. Consider: 3x3 + 1 = 10, which is EVEN. So the conjecture 3q + 1 always results in an odd number is simply NOT a true statement. You can show ten thousand examples of when 3q + 1 is an odd number, but even one case where 3q + 1 is an even number makes this conjecture false.